# Erasure-tolerance scheme for the surface codes on neutral atom quantum computers

### v3

@@ VII Discussion and Conclusion
In this paper, we have numerically verified the error correction performance of surface codes on systems where erasure errors continuously occur and accumulate on the array, such as in neutral atom quantum computers using optical tweezer arrays. Our work revealed that the surface code with longer code distances is more strongly affected by the accumulated erasure errors under the model where refilling erased qubits is impossible. The logical failure rate becomes badly high by increasing the code distance and repeating error correction because the threshold for the erasure error rate is theoretically zero. Additionally, iterating the syndrome measurements increases the number of erased qubits. These erased qubits shorten the effective code distance due to the formation of super stabilizers, which worsens the fault tolerance.
@@
In the following, we discuss aspects not considered in this study and open problems. The numerical calculations were performed under the assumption that the erasure error rate of each operation is uniform. However, neutral atom quantum computers are prone to leakage errors and atom loss, mainly when performing two-qubit gates [12, 30]. Therefore, investigating the effect of biased erasure error occurrence is a subject for future work.
@@
Investigating efficient scheduling and allocation of ancilla qubits for syndrome collection under this scheme should also be interesting. A compilation proposal that gradually changes the qubit allocations on the atom array has been proposed by [35]. By adopting such compilation of scheduling and routing approaches with syndrome measurement circuits, the Rydberg atoms should realize syndrome collection with more efficient circulation of ancilla qubits and more efficient scheduling of syndrome measurements than recent studies that assume only nearest-neighbor interactions [16]. Researching the scheduling with the ancilla-sharing strategy mentioned above will promote the feasible implementation of the super stabilizers under accumulated erasure errors.
@@
We could not analytically access the explicit effect of the accumulated erasure errors on decoding, which remains an open problem. The accumulated erasure errors can be modeled as temporally correlated, completely depolarizing errors. Evaluating these errors by mapping to the statistical mechanics model with long-range correlations [36] is expected to reveal the effect of accumulated erasure errors on the logical failure rate.

### v4

@@ VII Discussion and Conclusion
In this paper, we have numerically verified the error correction performance of surface codes on systems where erasure errors continuously occur and accumulate on the array, such as in neutral atom quantum computers using optical tweezer arrays. Our work revealed that the surface code with longer code distances is more strongly affected by the accumulated erasure errors under the model where refilling erased qubits is impossible. The logical failure rate becomes extremely high by increasing the code distance and repeating error correction because the threshold for the erasure error rate is theoretically zero. Additionally, iterating the syndrome measurements increases the number of erased qubits. These erased qubits shorten the effective code distance due to the formation of super stabilizers, which worsens the fault tolerance.
@@
In the following, we discuss aspects not considered in this study and open problems. The numerical calculations were performed under the assumption that the erasure error happens with the uniform rate before each syndrome measurement round. However, neutral atom quantum computers are prone to leakage errors and atom loss, mainly when performing two-qubit gates [13, 34]. Moreover, erased qubits should be detected with faulty components such that some leakage may spread before erasure conversion, while we assumed a noiseless erasure detection. Therefore, investigating the effect of biased erasure error occurrence at mid-syndrome measurement rounds and faulty erasure detections are a subject for future work.
@@
Investigating efficient scheduling and allocation of ancilla qubits for syndrome collection under this scheme should also be interesting. A compilation proposal that gradually changes the qubit allocations on the atom array has been proposed by [41]. By adopting such compilation of scheduling and routing approaches with syndrome measurement circuits, the Rydberg atoms should realize syndrome collection with more efficient circulation of ancilla qubits and more efficient scheduling of syndrome measurements than recent studies that assume only nearest-neighbor interactions [17]. Researching the scheduling with the ancilla-sharing strategy mentioned above will promote the feasible implementation of the super stabilizers under accumulated erasure errors. Moreover, a numerical simulation for repeating the transferring a logical qubit many times will reveal more detailed effect of $k$-shift erasure recovery for the logical error rate. However, we could not address this simulation because of the limitations of computational resources and runtime issues to perform such simulations. Providing the simulation of our framework on an efficient QEC simulation with erasure errors such as tensor-netowrk simulation[42] and Pauli+ [43] is a future work.
@@
We could not analytically derive the explicit effect of the accumulated erasure errors on decoding, which remains an open problem. The accumulated erasure errors can be modeled as temporally correlated, completely depolarizing errors. Evaluating these errors by mapping to the statistical mechanics model with long-range correlations [44] is expected to reveal the effect of accumulated erasure errors on the logical failure rate.